OpenStudy (anonymous):

Rewrite the expression as a trigonometric function of a single angle measure.

OpenStudy (anonymous):

\[\frac{ \tan 4\theta + \tan 2 \theta }{ 1 - \tan 4\theta + \tan 2 \theta }\]